OpenStudy (anonymous):

find all square roots. 4 answers x^4 - 8x^3 + 25x^2 = 0

OpenStudy (anonymous):

just factor it

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

ok what is your first step?

OpenStudy (anonymous):

all terms have an x, that means you can __________?

OpenStudy (anonymous):

factor out?

OpenStudy (anonymous):

can u use the quadratic formula in this equation

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

just answer my questions

OpenStudy (anonymous):

what is your first step? all terms have an x, that means you can factor out __________?

OpenStudy (anonymous):

the x^2

OpenStudy (anonymous):

ok so once you factor that out, what does the equation look like?

OpenStudy (anonymous):

x^2 ( x^2 - 8x + 25)

OpenStudy (anonymous):

= 0

OpenStudy (anonymous):

ok now remember the zero property of multiplication anythign multiplied by zero is equal to zero this means either x^2 is equal to zero or x^2-8x+25 is equal to zero now lets start by looking at x^2=0 what are the 2 roots (well one root) x=?

OpenStudy (anonymous):

in other words x^2 =0 x =?

OpenStudy (anonymous):

... 1^2= 1 1 is not equal to zero.....

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

2 of your roots would be zero x^2 = x *x =0 recalling zero property of multiplication x=0 and x=0

OpenStudy (anonymous):

do you understand this?

OpenStudy (anonymous):

yea so far.

OpenStudy (anonymous):

good, so we know two of the roots, now to find the other two recalling zero property of multiplication x^2-8x+25 must also be equal to zero x^2-8x+25=0 we need to factor this in order to determine the 2 roots unfortunately, if you try, you will realize that you cannot not factor it with normal means this is where the quadratic equation comes in

OpenStudy (anonymous):

in otherwords use the quadratic equation to determine hte other 2 roots

OpenStudy (anonymous):

okay thanks